OpenStudy (anonymous):

write a linear factorization of the function: f(x)=x^3+4x^2+25x+100

OpenStudy (jhannybean):

\[f(x) = x^3 +4x^2 +25x +100\]Split this function into 2 groups. \[=\color{blue}{(x^3+4x^2)} +\color{red}{(25x+100)}\]Now in each group, factor out the LCM. In this case it would be \(\color{blue}{x^2}\) from the first group and \(\color{red}{25}\) from the second group.\[=\color{red}{x^2}\color{blue}{(x+4)} +\color{red}{25}\color{blue}{(x+4)}\]You will see that once those are factored out, you have common likes terms as factors. You can combine these. \[f(x)=\color{red}{(x^2+25)}\color{blue}{(x+4)}\] You will see now that this is the linear factorized form of your function. \[f(x) = x^3 +4x^2 +25x+100 = (x^2+25)(x+4)\]

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

No problem :) And remember to medal the user that helped you the most! :-)